a biker travels a rate of 12 mphwhen there is no wind. Against the wind, he bikes 8 miles in the same time that it takes to bike 14 miles with the wind. what is the speed of the wind?
Answer by mananth(16946)

biker speed = 12 mph
windspeed = x mph

speed against wind 12 - x
speed with wind 12 + x

Distance against wind 8 miles
Distance with wind 14 miles

t=d/r time against = time with wind -
8/(12-x )=14/(12+x)8
8( 12 + x )= 14 (12 -x )
96 + 8 x = 168 - -14 x
14 x 8 x 168 + -96
22 x = 72
/ 22 check
x = 3.27 mph wind speed